OpenStudy (eric_d):

The radius of the Earth is 6400 km and it takes 24 hrs to make a complete rotation about it's axis.

OpenStudy (eric_d):

a)What is the angular velocity of the Earth's rotation about its axis in rev s^-1

OpenStudy (eric_d):

@Abhisar

OpenStudy (abhisar):

\(\sf \huge \omega =\frac{v}{r}\)

OpenStudy (abhisar):

\(\sf \huge v=\frac{2\pi r}{t}\)

OpenStudy (abhisar):

\(\sf \huge \omega = \frac{2\pi}{t}\)

OpenStudy (abhisar):

What's the answer ?

OpenStudy (eric_d):

The given answer is 1.16 x 10^-5 rev s^-1

OpenStudy (eric_d):

I'm getting 7.43 x 10^-5

OpenStudy (abhisar):

Lol ! It's easy one...we were complicating it :D

OpenStudy (abhisar):

Earth takes 24 hours to complete one revolution, i.e it takes 24*3600 seconds to complete its on erevolution. So angular velocity in revolution per second will be 1/24*3600 \(\approx 1.6\times 10^{-5}rev/s\)

OpenStudy (eric_d):

So, I don't hv to apply any formula here

OpenStudy (abhisar):

No....and radius is just given to fool u

OpenStudy (eric_d):

Let say the question mentions.. it takes 24 hrs to make 2 complete rotation, it'll be 2/86400

OpenStudy (abhisar):

yes...

OpenStudy (eric_d):

Ok.

OpenStudy (eric_d):

Thanks @Abhisar

OpenStudy (abhisar):

\(\rlap{\color{blue}{\huge\bigstar}\huge \color{blue}{ \text{You're Most Welcome! }}\color{blue}\bigstar}{\; \color{aqua}{\huge\bigstar}\huge \color{aqua}{\text{You're Most Welcome! }}\color{aqua}\bigstar}\) \(~~~~~~~~~~~~~~~~~~~~~~~~~~~\color{green}{\huge\ddot\smile}\color{blue}{\huge\ddot\smile}\color{pink}{\huge\ddot\smile}\color{red}{\huge\ddot\smile}\color{yellow}{\huge\ddot\smile}\)